OpenStudy (anonymous):

if sin theta=1/3, theta in quadrant II, find exact value of: b)sin(theta+pi/6)

OpenStudy (anonymous):

don't put 1/3 for a do I?

OpenStudy (anonymous):

sorry i dont get, i know about the formula, but not how to use it, theta?, what?+

OpenStudy (paxpolaris):

you can put in 1/3 for sinA

OpenStudy (paxpolaris):

B is pi/6

OpenStudy (paxpolaris):

all you need is to find cosA ... and you have everything you need to use the formula

OpenStudy (anonymous):

oh

OpenStudy (anonymous):

how

OpenStudy (anonymous):

and is pi/6 rad

OpenStudy (anonymous):

how wud i find exact value of 1/3

OpenStudy (anonymous):

so confused

OpenStudy (paxpolaris):

\[\large \sin \left( \theta+\frac \pi6 \right)=\sin(\theta)\cos \left( \frac \pi6 \right)+\cos(\theta)\sin\left( \frac \pi6 \right)\] using the formula we get this .. ^^^

OpenStudy (jhannybean):

Yes that :P

OpenStudy (paxpolaris):

yes pi/6 is in radians

OpenStudy (anonymous):

how do we use 1/3 for the formula, where is hat coming in

OpenStudy (jhannybean):

well, if \(\ \sin(\theta) = \frac{1}{3}\), then \(\ \theta = \sin^{-1}(\frac{1}{3})\)

OpenStudy (paxpolaris):

sin(θ) = 1/3 ... so , in the formula we can replace sin(θ) with 1/3.

OpenStudy (paxpolaris):

Hint: you don't need to find theta

OpenStudy (jhannybean):

I can't see your post because there is inumerous amount of blackened question marks x_x

OpenStudy (paxpolaris):

yeah , sorry ... it appears properly if you refresh.

OpenStudy (anonymous):

buts lets sin a = sin1/3 how wud i find exact value for thattt ?

OpenStudy (anonymous):

pi/6 i wud understand

OpenStudy (paxpolaris):

but the question does not say sin a = sin 1/3 ...???

OpenStudy (anonymous):

no theta = 1/3

OpenStudy (anonymous):

so how do i find a??

OpenStudy (paxpolaris):

sin theta = 1/3 there is no need to find theta.... we only need to find sin(theta + 6)

OpenStudy (jhannybean):

Ohh! yeah pax is right , haha. they're telling you that instead of finding the theta value for sin( theta), to replace sin(theta) with 1/3. When you use the formula Pax and Akash posted, you will find the exact value you are looking for :)

OpenStudy (anonymous):

ehh

OpenStudy (paxpolaris):

\[\sin(\theta)=1/3\ ...\{\because given\}\]\[\sin\left( \frac \pi6\right)=\frac12\]\[\cos\left( \frac \pi6\right)=\frac{\sqrt3}2\] plug these numbers in \[\large \sin \left( \theta+\frac \pi6 \right)=\sin(\theta)\cos \left( \frac \pi6 \right)+\cos(\theta)\sin\left( \frac \pi6 \right)\]\[\large \implies \sin \left( \theta+\frac \pi6 \right)=\frac13 \cdot {\sqrt3 \over 2}+\color{blue}{\cos(\theta)}\cdot \frac12\]

OpenStudy (paxpolaris):

so we have all the numbers in the formula except \(\color{blue}{\cos(\theta)}\) ... which we can find using pythagorean theorem.

OpenStudy (anonymous):

got it

OpenStudy (anonymous):

vos wud be -2sqrt2 over 3

OpenStudy (anonymous):

plug in

OpenStudy (anonymous):

and u get sqrt 3 -2sqrt2 over6

OpenStudy (anonymous):

i overthink

OpenStudy (anonymous):

:p

OpenStudy (jhannybean):

wait wait.. how would we use the pythagorean theorem to find cos?

OpenStudy (anonymous):

sin=o/h

OpenStudy (paxpolaris):